Sorry you are right I did not type it in correctly. I do need to solve it in quadratic form
X^2/5-x^1/5-2=0
Sorry!

Answer by stanbon(75887)

Sorry you are right I did not type it in correctly. I do need to solve it in quadratic form
X^2/5-x^1/5-2=0
----------------
Factor:
(x^(1/5)-2)(x^(1/5)+1) = 0
---
x^(1/5) = 2 or x^(1/5) = -1
x = 32 or x = -1
------------------------
Cheers,
Stan H.
==========